# Can Large Language Models Adequately Perform Symbolic Reasoning Over Time Series?

### v3

@@ Abstract
Uncovering hidden symbolic laws from time series data, as an aspiration dating back to Kepler’s discovery of planetary motion, remains a core challenge in scientific discovery and artificial intelligence. While Large Language Models show promise in structured reasoning tasks, their ability to infer interpretable, context-aligned symbolic structures from time series data is still underexplored. To systematically evaluate this capability, we introduce SymbolBench [^1] https://anonymous.4open.science/r/SymbolBench-6F40, a comprehensive benchmark designed to assess symbolic reasoning over real-world time series across three tasks: multivariate symbolic regression, Boolean network inference, and causal discovery. Unlike prior efforts limited to simple algebraic equations, SymbolBench spans a diverse set of symbolic forms with varying complexity. We further propose a unified framework that integrates LLMs with genetic programming to form a closed-loop symbolic reasoning system. Our empirical results reveal key strengths and limitations of current models, highlighting the importance of combining domain knowledge, context alignment, and reasoning structure to improve LLMs in automated scientific discovery.
Can Large Language Models Adequately Perform Symbolic Reasoning Over Time Series?
Zewen Liu1, Juntong Ni1, Xianfeng Tang2, Max S.Y. Lau1, Wenpeng Yin3, Wei Jin1 1Emory University, 2Amazon, 3Penn State University {zewen.liu, juntong.ni, msy.lau, wei.jin}@emory.edu, xianfengtang@outlook.com; wenpeng@psu.edu
@@ 5 Experiment
Table 2: Symbolic regression results for CDEs across 4 dimensions. We use percentage for $SR^{2}$ and $ACC_{0.9}$. Yellow, Orange, and Cyan mark the first, second, and third place, respectively. Statistically significant improvements over the best baseline (paired t-test, $p$ < 0.05) are underlined.
@@
Pro. PySR ODE. Naïve Base Ctx CoT Ctx CoT Ctx CoT Ctx CoT Ctx CoT Ctx CoT
Dim=1
Complexity $\downarrow$ $5.06$ $2.91$ $4.68$ $4.17$ $5.14$ $5.34$ $3.91$ $1.89$ $1.97$ $2.63$ $2.69$ $2.31$ $2.36$ $2.37$ $2.09$ $3.29$ $3.00$
Dim=2
Complexity $\downarrow$ $5.71$ $7.00$ $9.05$ $8.76$ $9.75$ $10.00$ $9.07$ $6.93$ $7.07$ $7.93$ $8.35$ $6.91$ $6.64$ $6.33$ $6.33$ $9.88$ $9.63$
Dim=3
Complexity $\downarrow$ $7.87$ $8.82$ $13.90$ $16.60$ $17.40$ $18.30$ $15.50$ $15.70$ $17.30$ $16.30$ $15.40$ $12.60$ $12.30$ $12.00$ $11.50$ $17.00$ $17.20$
Dim=4
Complexity $\downarrow$ $9.70$ $10.60$ $14.40$ $22.90$ $23.20$ $23.10$ $18.50$ $20.60$ $19.60$ $17.90$ $18.70$ $14.50$ $13.90$ $13.90$ $13.60$ $22.60$ $22.10$
@@ 5.2 Experimental Results
Across all three tasks, we select LLMs with various sizes and architectures, including Qwen2.5-14B Team (2024), Llama-3.2B Dubey et al. (2024), Mathstral-7B Jiang et al. (2023), GPT-4o-mini Achiam et al. (2023), and ChatTS-14B Xie et al. (2024). For GPT-4o-mini and ChatTS-14B, we use inputs with visual and temporal modalities. We present the results in Table 2,  3, and  4, where ablation studies on the four prompts with Qwen2.5-14B are integrated into the main tables rather than reported in a dedicated section.
@@
Obs. 2: LLMs’ performance degrades with problem difficulty. Across all models, performance consistently declines as the dimensionality of the system increases. While conventional methods such as PySR remain competitive in low-dimensional systems (e.g., dim = 1), baseline models tend to exhibit a steeper performance drop compared to LLMs. Similarly, there is a notable drop in accuracy when models are evaluated on OOD data, underscoring the increased complexity and generalization challenges posed by these scenarios.
Obs. 3: Chain-of-thought prompting does not consistently improve performance. Except ChatTS-14B, further introducing CoT prompting does not lead to consistent gains, especially on the CDEs dataset. While increasing test-time compute has been shown to improve outcomes in various tasks Snell et al. (2024), to which CoT can contribute, similar limitations are observed in other benchmarks like SciBench Wang et al. (2023).
@@
Table 3: Comparison of Boolean network inference across ID and OOD settings. Full results in Appendix E.
@@ Appendix A Related Work
LLM Symbolic Reasoning. LLMs exhibit strong in-context learning and reasoning capabilities, enabling them to perform tasks such as logical inference Ahn et al. (2024); Wang and Chen (2023) and even preliminary forms of temporal reasoning Fang et al. (2024). Prior work on time series reasoning with LLMs, however, has largely emphasized linguistic outputs, for example, generating natural-language summaries of temporal patterns, identifying anomalies, or extracting descriptive insights from sequential data. While valuable, these approaches focus on surface-level interpretation and do not attempt to uncover the underlying symbolic structures that govern dynamic systems. In contrast, our work targets symbolic reasoning tasks where the output is not a narrative but a formal structure: scientific equations, logical rules, or causal graphs. These forms require abstraction, generalization, and alignment with domain context, going beyond function generation. Existing efforts to apply LLMs to symbolic regression have mainly treated them as equation generators Merler et al. (2024); Li et al. (2024); Shojaee et al. (2024), producing functions that may fit data but often lack explanatory power or contextual plausibility. Moreover, the symbolic space considered so far has been restricted almost exclusively to algebraic expressions, leaving other important forms such as Boolean networks Zhang et al. (2024) and causal relations Assaad et al. (2022) largely unexplored.
@@ Appendix D Baseline Implementation
Appendix E Full Results on the BN and SCM datasets
@@ Appendix I Convergence Rate
(b) BN
(c) SCM
Figure 9: Convergence Rate on CDE, BN, and SCM datasets.

### v4

# SymbolBench: Can Large Language Models Adequately Perform Symbolic Reasoning Over Time Series?

@@ Abstract
Uncovering hidden symbolic laws from time series data, as an aspiration dating back to Kepler’s discovery of planetary motion, remains a core challenge in scientific discovery and artificial intelligence. While Large Language Models show promise in structured reasoning tasks, their ability to infer interpretable, context-aligned symbolic structures from time series data is still underexplored. To systematically evaluate this capability, we introduce SymbolBench, a comprehensive benchmark designed to assess symbolic reasoning over real-world time series across three tasks: multivariate symbolic regression, Boolean network inference, and causal discovery. Unlike prior efforts limited to simple algebraic equations, SymbolBench spans a diverse set of symbolic forms with varying complexity. We further propose a unified framework that integrates LLMs with genetic programming to form a closed-loop symbolic reasoning system. Our empirical results reveal key strengths and limitations of current models, highlighting the importance of combining domain knowledge, context alignment, and reasoning structure to improve LLMs in automated scientific discovery. https://github.com/nuuuh/SymbolBench
SymbolBench: Can Large Language Models Adequately Perform Symbolic Reasoning Over Time Series?
Zewen Liu1, Juntong Ni1, Xianfeng Tang2, Max S.Y. Lau1, Qi He3, Wenpeng Yin4, Wei Jin1, 1Emory University, 2Amazon, 3Microsoft, 4Penn State University {zewen.liu, juntong.ni, msy.lau, wei.jin}@emory.edu, qhe@microsoft.com; xianfengtang@outlook.com; wenpeng@psu.edu
@@ 5 Experiment
Table 2: Symbolic regression results for CDEs across 4 dimensions. We use percentage for $SR^{2}$ and $ACC_{0.9}$. Orange and Yellow mark the first and second place, respectively. Statistically significant improvements over the best baseline (paired t-test, $p$ < 0.05) are underlined.
@@
Pro. PySR ODE. Ctx CoT Ctx CoT Ctx CoT Ctx CoT Ctx CoT Ctx CoT
Dim=1 Complexity $\downarrow$ $5.06$ $2.91$ $4.68$ $5.34$ $3.91$ $1.89$ $1.97$ $2.63$ $2.69$ $2.31$ $2.36$ $2.37$ $2.09$ $3.29$ $3.00$
Symbolic Prox. $\downarrow$ $4.86$ $3.54$ $5.18$ $5.57$ $4.89$ $4.03$ $4.14$ $3.86$ $4.09$ $3.89$ $3.79$ $3.97$ $4.06$ $4.83$ $4.37$
ID $SR^{2}$ $95.50$ $96.90$ $83.40$ 99.00 97.40 $95.10$ $94.90$ $96.20$ $96.20$ $92.80$ $95.40$ $95.70$ $95.30$ $95.10$ $97.00$
$ACC_{0.9}$ $91.40$ $97.10$ $73.50$ 100.00 $97.10$ $91.40$ $91.40$ $97.10$ $97.10$ $94.30$ $97.00$ $97.10$ $97.10$ $91.40$ $97.10$
OOD $SR^{2}$ $74.10$ $88.20$ $54.10$ $61.80$ $61.80$ $69.90$ $64.60$ $71.00$ $66.90$ $70.80$ $66.20$ $71.50$ $62.90$ $63.80$ $70.30$
$ACC_{0.9}$ $65.70$ $85.70$ $47.10$ $52.90$ $51.40$ $51.40$ $48.60$ $62.90$ $60.00$ $54.30$ $51.50$ $54.30$ $45.70$ $51.40$ $54.30$
Avg. ID Rank $\downarrow$ $10.5$ $3.5$ $17.0$ $1.0$ $2.0$ $12.0$ $13.0$ $4.0$ $4.0$ $13.5$ $10.0$ $5.0$ $6.5$ $12.0$ $3.0$
Avg. OOD Rank $\downarrow$ $2.0$ $1.0$ $16.0$ $11.5$ $12.5$ $9.5$ $12.5$ $4.5$ 77 66 $10$ 55 $14.5$ $11.5$ $6.5$
Dim=4 Complexity $\downarrow$ $9.70$ $10.60$ $14.40$ $23.10$ $18.50$ $20.60$ $19.60$ $17.90$ $18.70$ $14.50$ $13.90$ $13.90$ $13.60$ $22.60$ $22.10$
Symbolic Prox. $\downarrow$ $32.70$ $30.30$ $36.40$ $34.10$ $32.90$ $34.70$ $32.70$ $32.50$ $32.40$ $34.80$ $32.90$ $32.20$ $31.90$ $38.00$ $36.30$
ID $SR^{2}$ $11.50$ $74.60$ $34.80$ 93.30 91.00 $49.40$ $61.10$ $78.00$ $78.70$ $88.60$ $86.90$ $86.10$ $89.10$ $78.70$ $80.00$
$ACC_{0.9}$ $8.70$ $65.20$ $29.60$ 93.90 89.30 $40.70$ $52.00$ $74.10$ $75.00$ $78.80$ $70.00$ $71.40$ $85.70$ $74.10$ $80.80$
OOD $SR^{2}$ $7.00$ $23.80$ $11.40$ $19.40$ 34.60 $28.30$ $21.30$ 30.70 36.90 26.90 27.60 $30.40$ $30.50$ $25.00$ $24.50$
$ACC_{0.9}$ $0.00$ $21.70$ $7.10$ $12.50$ 33.30 $12.00$ $13.60$ $24.00$ $30.80$ $19.40$ $20.70$ $15.40$ $21.40$ $21.70$ $17.40$
Avg. ID Rank $\downarrow$ $17.0$ $13.0$ $16.0$ $2.0$ $3.5$ $15.0$ $14.0$ $10.5$ $9.0$ $6.5$ $9.5$ $9.5$ $5.0$ $9.5$ $7.5$
Avg. OOD Rank $\downarrow$ $17.0$ $8.0$ $16.0$ $14.0$ $1.5$ $11.0$ $13.0$ $3.0$ $1.5$ $8.5$ $7.5$ $8.5$ $5.0$ $7.0$ $10.5$
Figure 4: Performance improvement of varying prompting strategies (Base, CTX, CoT) over the Naïve baseline. The context-aware strategy (CTX) consistently achieves the largest gains across metrics.
@@ 5.2 Experimental Results
Across all three tasks, we select LLMs with various sizes and architectures, including Qwen2.5-14B Team (2024), Llama-3.2-3B Dubey et al. (2024), Mathstral-7B Jiang et al. (2023), GPT-4o-mini Achiam et al. (2023), and ChatTS-14B Xie et al. (2024). For GPT-4o-mini and ChatTS-14B, we use inputs with visual and temporal modalities. We present the results in Table 2,  3, and  4, where ablation studies on the four prompts with Qwen2.5-14B are integrated into the main tables rather than reported in a dedicated section.
@@
Obs. 2: LLMs’ performance degrades with problem difficulty. Across all models, performance consistently declines as the dimensionality of the system increases. While conventional methods such as PySR remain competitive in low-dimensional systems (e.g., dim = 1) in Table 2, baseline models tend to exhibit a steeper performance drop compared to LLMs on higher dimensional systems (full results in Appendix E). Similarly, there is a notable drop in accuracy when models are evaluated on OOD data, underscoring the increased complexity and generalization challenges posed by these scenarios.
Obs. 3: Chain-of-thought prompting does not consistently improve performance. Except ChatTS-14B, further introducing CoT prompting does not lead to consistent gains, especially on the CDEs dataset, as shown in Figure 4. While increasing test-time compute has been shown to improve outcomes in various tasks Snell et al. (2024), to which CoT can contribute, similar limitations are observed in other benchmarks, e.g., SciBench Wang et al. (2023b).
@@
Table 3: Comparison of Boolean network inference across ID and OOD settings. Full results in Appendix E. Orange and Yellow mark the first and second place, respectively.
@@ 7 Limitations
Acknowledgments
Zewen Liu and Wei Jin are supported by the U.S. National Science Foundation under Award Number 2504088, and the National Institute of Allergy and Infectious Diseases of the NIH under Award Numbers R01AI189829 and R01AI197111. The content is the sole responsibility of the authors and does not necessarily represent the views of the NIH. This project is also partially supported by the NAIRR Pilot Program.
@@ Appendix A Related Work
LLM Symbolic Reasoning. LLMs exhibit strong in-context learning and reasoning capabilities, enabling them to perform tasks such as logical inference Ahn et al. (2024); Wang and Chen (2023) and even preliminary forms of temporal reasoning Fang et al. (2024). Recent LLM systems also benefit from combining structured and textual context Glenn et al. (2024); Lei et al. (2025); Liu et al. (2025); Wang et al. (2023a). Prior work on time series reasoning with LLMs, however, has largely emphasized linguistic outputs, for example, generating natural-language summaries of temporal patterns, identifying anomalies, or extracting descriptive insights from sequential data. While valuable, these approaches focus on surface-level interpretation and do not attempt to uncover the underlying symbolic structures that govern dynamic systems. In contrast, our work targets symbolic reasoning tasks where the output is not a narrative but a formal structure: scientific equations, logical rules, or causal graphs. These forms require abstraction, generalization, and alignment with domain context, going beyond function generation. Existing efforts to apply LLMs to symbolic regression have mainly treated them as equation generators Merler et al. (2024); Li et al. (2024); Shojaee et al. (2024), producing functions that may fit data but often lack explanatory power or contextual plausibility. Moreover, the symbolic space considered so far has been restricted almost exclusively to algebraic expressions, leaving other important forms such as Boolean networks Zhang et al. (2024) and causal relations Assaad et al. (2022) largely unexplored.
@@ Appendix D Baseline Implementation
Appendix E Full Results
We present full results across all dimensions in Table 9, and the performance of more LLMs on BN and SCM discovery in Table 10 and Table 11.
Table 9: Symbolic regression results for CDEs across 4 dimensions. We use percentage for $SR^{2}$ and $ACC_{0.9}$. Orange,Yellow, and Cyan mark the first, second, and third place, respectively. Statistically significant improvements over the best baseline (paired t-test, $p$ < 0.05) are underlined.
Pro. PySR ODE. Naïve Base Ctx CoT Ctx CoT Ctx CoT Ctx CoT Ctx CoT Ctx CoT
Dim=1 Complexity $\downarrow$ $5.06$ $2.91$ $4.68$ $4.17$ $5.14$ $5.34$ $3.91$ $1.89$ $1.97$ $2.63$ $2.69$ $2.31$ $2.36$ $2.37$ $2.09$ $3.29$ $3.00$
Dim=2 Complexity $\downarrow$ $5.71$ $7.00$ $9.05$ $8.76$ $9.75$ $10.00$ $9.07$ $6.93$ $7.07$ $7.93$ $8.35$ $6.91$ $6.64$ $6.33$ $6.33$ $9.88$ $9.63$
Dim=3 Complexity $\downarrow$ $7.87$ $8.82$ $13.90$ $16.60$ $17.40$ $18.30$ $15.50$ $15.70$ $17.30$ $16.30$ $15.40$ $12.60$ $12.30$ $12.00$ $11.50$ $17.00$ $17.20$
Dim=4 Complexity $\downarrow$ $9.70$ $10.60$ $14.40$ $22.90$ $23.20$ $23.10$ $18.50$ $20.60$ $19.60$ $17.90$ $18.70$ $14.50$ $13.90$ $13.90$ $13.60$ $22.60$ $22.10$
@@ Appendix I Convergence Rate
Figure 10: Convergence Rate on the CDE dataset.